# Intrinsic Second-Order magnon Thermal Hall Effect

### v1

@@ Appendix D The derivation of the heat current density correction induced by TBCP
From Eq. (80) and keeping to the second order of gradient of temperature, we get the intrinsic second-order thermal Hall current of magnon as
@@ Appendix E Further details of model
$\begin{split}\hat{H}&=\sum_{\langle ij\rangle}J_{ij}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}+\sum_{\ll ij\gg}\mathbf{D}_{ij}\cdot\left(\hat{\mathbf{S}}_{i}\times\hat{\mathbf{S}}_{j}\right)+g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}(\mathbf{r})+K\sum_{i}\hat{S}_{i}^{z2}\\ \end{split}$ (82)
We divide the Hamiltonian into three parts, $\hat{H}_{1}=\sum_{\langle ij\rangle}J_{ij}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}$, $\hat{H}_{2}=\sum_{\ll ij\gg}\mathbf{D}_{ij}\cdot\left(\hat{\mathbf{S}}_{i}\times\hat{\mathbf{S}}_{j}\right)$, $\hat{H}_{3}=g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}(\mathbf{r})$, and $\hat{H}_{4}=K\sum_{i}\hat{S}_{i}^{z2}$.
@@
$\begin{split}\hat{H}_{4}=K\sum_{i}(S^{2}-2S\hat{a}_{i}^{{\dagger}}\hat{a}_{i})+K\sum_{i}(S^{2}-2S\hat{b}_{i}^{{\dagger}}\hat{b}_{i}).\end{split}$ (88)
@@
$\begin{split}\hat{H}_{4}&=-2KS\sum_{\mathbf{k}}(\hat{a}_{\mathbf{k}}^{{\dagger}}\hat{a}_{\mathbf{k}}+\hat{b}_{\mathbf{k}}^{{\dagger}}\hat{b}_{\mathbf{k}}),\end{split}$ (92)

### v2

@@ Appendix D The derivation of the heat current density correction induced by TBCP
From Eq. (81) and keeping to the second order of gradient of temperature, we get the intrinsic second-order Hall heat current of magnon as
@@ Appendix E Further details of model
$\begin{split}\hat{H}&=\sum_{\langle ij\rangle}J_{ij}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}+\sum_{\ll ij\gg}\mathbf{D}_{ij}\cdot\left(\hat{\mathbf{S}}_{i}\times\hat{\mathbf{S}}_{j}\right)+g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}(\mathbf{r})+\mathcal{K}\sum_{i}\hat{S}_{i}^{z2}\\ \end{split}$ (83)
We divide the Hamiltonian into three parts, $\hat{H}_{1}=\sum_{\langle ij\rangle}J_{ij}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}$, $\hat{H}_{2}=\sum_{\ll ij\gg}\mathbf{D}_{ij}\cdot\left(\hat{\mathbf{S}}_{i}\times\hat{\mathbf{S}}_{j}\right)$, $\hat{H}_{3}=g_{J}\mu_{B}\sum_{i}\hat{\mathbf{S}}_{i}\cdot\mathbf{B}(\mathbf{r})$, and $\hat{H}_{4}=\mathcal{K}\sum_{i}\hat{S}_{i}^{z2}$.
@@
$\begin{split}\hat{H}_{4}=\mathcal{K}\sum_{i}(S^{2}-2S\hat{a}_{i}^{{\dagger}}\hat{a}_{i})+\mathcal{K}\sum_{i}(S^{2}-2S\hat{b}_{i}^{{\dagger}}\hat{b}_{i}).\end{split}$ (89)
@@
$\begin{split}\hat{H}_{4}&=-2\mathcal{K}S\sum_{\mathbf{k}}(\hat{a}_{\mathbf{k}}^{{\dagger}}\hat{a}_{\mathbf{k}}+\hat{b}_{\mathbf{k}}^{{\dagger}}\hat{b}_{\mathbf{k}}),\end{split}$ (93)